# Critical Phase Transition in a Large Language Model

### v1

@@ 6 Limitations
  • Our work is solely based on numerical analysis, and we have not provided any proof for the existence of the phase transition mathematically.
  • We only focus on POS sequences. Therefore, statistical quantities in this research do not capture semantic structures.
Acknowledgments and Disclosure of Funding
@@ Appendix G Critical Properties of WikiText
Appendix H Analysis of japanese-gpt2-medium
In the main text, we have investigated statistical properties of POS sequences generated by GPT-2 with varying temperature and have found the phase transition between repetitive sequences and incomprehensible ones. An interesting question is how universal this phase transition is. Here, we analyze japanese-gpt2-medium [Zhao and Sawada, , Sawada et al., 2024] using ja_core_news_sm pipeline from spaCy library [Montani et al., 2023]. For each $T$ and $N$, $2.4\times 10^{3}$ sequences were sampled.
@@ Appendix H Analysis of japanese-gpt2-medium
Figure 24: (A) Integrated correlation $\tau=\tau_{\text{PROPN},{\text{PROPN}}}$ in sequences generated by japanese-gpt2-medium as a function of temperature $T$ for various sequence length $N$. (B) The same quantity as a function of sequence length $N$ for various temperatures $T$. The black line represents a line proportional to $N$.

### v2

@@ 7 Limitations
  • The size of the models we analyzed is moderate (70M, 124M, and 361M). As a future direction, statistical analysis of larger models is important, because the true phase transition seems to exist in the large context window limit as we have discussed.
  • Since different natural languages widely exhibit critical behaviors, the criticality should be a necessary condition for generated texts to be natural. Further investigation is needed to better understand how the criticality relates to linguistic properties and practical tasks.
Acknowledgments
@@ Appendix G Critical Properties of WikiText
Appendix H Analysis of Japanese GPT-2
In the main text, we have investigated statistical properties of POS sequences generated by GPT-2 with varying temperature and have found the phase transition between repetitive sequences and incomprehensible ones. An interesting question is if this phase transition occur with other languages and LLMs. Here, we analyze Japanese GPT-2 medium with 361M parameters [Zhao and Sawada, , Sawada et al., 2024] using ja_core_news_sm pipeline from spaCy library [Montani et al., 2023]. For each $T$ and $N$, $2.4\times 10^{3}$ sequences were sampled. Error bars in the figures represent an $80\%$ confidence interval estimated using the symmetric bootstrap-$t$ method [Hall, 1988].
@@ Appendix H Analysis of Japanese GPT-2
Figure 24: (A) Integrated correlation $\tau=\tau_{\text{PROPN},{\text{PROPN}}}$ in sequences generated by Japanese GPT-2 as a function of temperature $T$ for various sequence length $N$. (B) The same quantity as a function of sequence length $N$ for various temperatures $T$. The black line represents a line proportional to $N$.
@@
Appendix I Analysis of Pythia
In the main text, we have generated texts with GPT-2 and mapped them to POS sequences, to show the existence of a phase transition. In this appendix, to examine if the phase transition is observed with other models and mappings, we use Pythia with 70M parameters [Biderman et al., 2023] and a character-based mapping. First, we transformed texts to POS sequences in the same manner as in the main text. Second, we also transformed texts to sequences of numerals, where “A” or “a” is mapped to 0, “B” or “b” is mapped to 1, … , “Z” or “z” is mapped to 25, other characters to 26, and space, tab, etc. to 27. For each $T$ and $N$, $10^{4}$ sequences were sampled. Error bars in the figures represent an $80\%$ confidence interval estimated using the symmetric bootstrap-$t$ method [Hall, 1988].
As shown in Fig. 29, the integrated correlation between PROPN and PROPN shows a behavior similar to that in the main text, suggesting that it diverges at low temperatures and converges at high temperatures. Meanwhile, Fig. 30 shows the integrated correlation between 26 and 26 in character-based sequences, which has the largest contribution. The result is also similar to that for GPT-2. These observations support that the phase transition occurs with different LLMs and mappings.
Figure 29: (A) Integrated correlation $\tau=\tau_{\text{PROPN},\text{PROPN}}$ in POS sequences generated by Pythia as a function of temperature $T$. (B) The same quantity as a function of sequence length $N$. The black line represents a line proportional to $N$.
Figure 30: (A) Integrated correlation between 26 (corresponding to other characters) and 26 in numeral sequences obtained by the character-based mapping from texts generated by Pythia, as a function of temperature $T$ (B) The same quantity as a function of sequence length $N$. The black line represents a line proportional to $N$.